The regression equation is y = 47.17 - 9.74 x1 + 0.428 x2 + 18.2 x3. If x1=14.5, x2=220, x3=5, the predicted value of y is? A) 373.56 B) 91.1 C) 47.17 D) 0
Answer #1

The regression equation is:

y = 47.17 - (9.74)x1 + (0.428)x2 + (18.2)x3

Putting the values, x1=14.5, x2=220 and x3=5, we get,

y = 47.17 - (9.74)(14.5) + (0.428)(220) + (18.2)(5) = 91.1

Hence option B is the correct option.